# A Unified Framework for Analyzing Meta-algorithms in Online Convex Optimization

### v3

@@
Abstract
In this paper, we analyze the problem of online convex optimization in different settings, including different feedback types (full-information/semi-bandit/bandit/etc) in either stochastic or non-stochastic setting and different notions of regret (static adversarial regret/dynamic regret/adaptive regret). This is done through a framework which allows us to systematically propose and analyze meta-algorithms for the various settings described above. We show that any algorithm for online linear optimization with fully adaptive adversaries is an algorithm for online convex optimization. We also show that any such algorithm that requires full-information feedback may be transformed to an algorithm with semi-bandit feedback with comparable regret bound. We further show that algorithms that are designed for fully adaptive adversaries using deterministic semi-bandit feedback can obtain similar bounds using only stochastic semi-bandit feedback when facing oblivious adversaries. We use this to describe general meta-algorithms to convert first order algorithms to zeroth order algorithms with comparable regret bounds. Our framework allows us to analyze online optimization in various settings, recovers several results in the literature with a simplified proof technique, and provides new results.
1 Introduction
1.1 Overview
Online optimization problems represent a class of problems where the decision-maker, referred to as an agent, aims to make sequential decisions in the face of an adversary with incomplete information (?, ?). This setting mirrors a repeated game, where the agent and the adversary engage in a strategic interplay over a finite time horizon, commonly denoted as $T$. The dynamics of this game involve the agent’s decision-making process, the adversary’s strategy, and the information exchange through a query oracle.
@@ 1.1 Overview
One of our key contribution lies in establishing that any semi-bandit feedback online linear (or quadratic) optimization algorithm for fully adaptive adversary is also an online convex (or strongly convex) optimization algorithm. While the above result is for semi-bandit feedback, we then show that in online convex optimization for fully adaptive adversary, semi-bandit feedback is generally enough and more information is not needed. Further, we show that algorithms that are designed for fully adaptive adversaries using deterministic semi-bandit feedback can obtain similar bounds using only stochastic semi-bandit feedback when facing oblivious adversaries. Finally, we introduce meta algorithms, based on the variant of classical result of (?, ?), that convert algorithms for semi-bandit feedback, such as the ones mentioned above, to algorithms for bandit or zeroth-order feedback. In addition to recovering many results in the literature with simplified proofs, we give new static regret guarantees for online strongly convex optimization and adaptive regret guarantees for convex optimization with deterministic zeroth order feedback.
Figure 1: Summary of the main results
@@ 1.2 Contributions
  • (I)
    We provide a framework that simplifies analysis of meta-algorithms in online learning.
  • (II)
    Using this framework, we formulate several meta-algorithms that facilitate the conversion of algorithms across various settings.
  • (III)
    In particular, we convert several insights in the literature into precise theorems, for example the idea of using 1-point gradient estimator in online optimization is formalized in Meta-algorithms 2 and 3 and Theorems 6 and 7. Similarly, the insight that “knowing the value of each loss function at two points is almost as useful as knowing the value of each function everywhere” (See (?)) is formalized in Meta-algorithm 4 and Theorem 8.
@@
(i) The first key result in this paper is Theorem 2. This result shows that any algorithm for online linear optimization with stochastic subgradient feedback results in a comparable regret with if the function class is replaced by a class of convex functions. This is the key result that allows us to go from online linear optimization to convex optimization in Figure 1 with semi-bandit feedback. We note that this result recovers the result for online convex game in (?), where the authors analyzed online convex optimization with deterministic gradient feedback and fully adaptive adversary and showed that the min-max regret bounds of online convex optimization is equal to that of online linear optimization (or online quadratic optimization is all convex functions are $\mu$-strongly convex).
@@
(iii) We note that the above results assume that the feedback oracle is deterministic. In the next key result, given in Theorem 5, we relax this assumption. We show that any online optimization algorithm for fully adaptive adversaries using deterministic semi-bandit feedback obtains similar regret bounds when facing an oblivious adversary using only stochastic semi-bandit feedback. Thus, this result allows us to get the results for online convex optimization with stochastic semi-bandit feedback from the results with deterministic semi-bandit feedback, as can be seen in Fig. 1. This result gives Theorem 6.5 in (?) as a special case. Further, this allows us to have a result that generalizes both adversarial regret and stochastic regret, as they are commonly defined in the literature.
@@
(iv) The above discussion was mostly focused on full information and semi-bandit feedback settings. However, in many practical setups, we only have access to zeroth order or bandit feedback. In the next key result, given in Theorems 6-7, we use the smoothing trick (?, ?, ?), in order to estimate the gradient from the value oracle. Using this estimated gradient, the results for zeroth order and bandit feedback are derived from that for first order and semi-bandit feedback, respectively. These results apply to any optimization problem, not necessarily convex. This approach recovers the results in (?) with the base algorithm of Online Gradient Descent ($\mathtt{OGD}$). In particular, this approach also matches the SOTA dynamic and adaptive regret bounds for online convex optimization with bandit feedback (?, ?, ?) with a simplified proof.
@@
(v) In the above results, we assumed that we only have access to an unbiased estimate of the zeroth order (or bandit) feedback. In (?), the setting of multi-point bandit feedback is introduced. Even though their setting and notion of regret is different from our work, one insight is relevant. Their analysis suggests that, in some cases, if we have access to exact values of the functions, knowing the value of each loss function at two points is almost as useful as knowing the value of each function everywhere. We formalize this insight by designing a meta-algorithm based on the two-point gradient estimator (?, ?) that allows us to convert algorithms for stochastic first order feedback into algorithms for deterministic zeroth order feedback with the same order of regret bound (Theorem 8). Note that this result holds for convex and non-convex optimization as it only relies on the function class being Lipschitz. This result is shown to recover the result for dynamic regret for Improved Ader (?), while achieving two new results for adaptive regret for convex function of $O(\sqrt{T})$ and static regret for strongly convex function of $O(\log T)$.
@@
Our work sheds light on the relation between different problems, algorithms and techniques used in online optimization and provides general meta-algorithms that allow conversion between different problem settings. For instance, we can use the results for a deterministic algorithm for online linear optimization with oblivious adversaries and deterministic semi-bandit feedback to that for a stochastic algorithm for online convex optimization with oblivious adversaries and stochastic bandit feedback using the different arrows as in Fig. 1. The key technical novelty is the recognition that many results, previously understood only in very limited contexts, actually apply more broadly. This broader applicability not only simplifies the analysis of numerous existing results within the field but also enables the derivation of multiple new results. For instance, results in (?, ?, ?, ?, ?, ?) as mentioned above can be demonstrated as corollaries of these more generalized results. Further, we provide new results, which shows that adaptive regret for convex function is $O(\sqrt{T})$ and static regret for strongly convex function is $O(\log T)$.
1.3 Related works
@@
The term “online convex optimization” was first defined in (?), where Online Gradient Descent ($\mathtt{OGD}$) was proposed, while this setting was first introduced in (?). The Follow-The-Perturbed-Leader ($\mathtt{FTPL}$) for online linear optimization was introduced by (?), inspired by Follow-The-Leader algorithm originally suggested by (?). Follow-The-Regularized-Leader ($\mathtt{FTRL}$) in the context of online convex optimization was coined in (?, ?). The equivalence of $\mathtt{FTRL}$ and Online Mirror Descent ($\mathtt{OMD}$) was observed by (?). Non-stationary measures of regret have been studied in  (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?) and many other papers. We refer the reader to surveys (?) and (?) for more details.
2 Background and Notations
@@
For a set ${\mathcal{D}}\subseteq{\mathbb{R}}^{d}$, we define its affine hull $\operatorname{aff}({\mathcal{D}})$ to be the set of $\alpha{\mathbf{x}}+(1-\alpha){\mathbf{y}}$ for all ${\mathbf{x}},{\mathbf{y}}$ in ${\mathcal{D}}$ and $\alpha\in{\mathbb{R}}$. The relative interior of ${\mathcal{D}}$ is defined as
$\operatorname{relint}({\mathcal{D}}):=\{{\mathbf{x}}\in{\mathcal{D}}\mid\exists r>0,{\mathbb{B}}_{r}({\mathbf{x}})\cap\operatorname{aff}({\mathcal{D}})\subseteq{\mathcal{D}}\}.$
@@
We use ${\mathbb{B}}_{r}({\mathcal{D}})$ to denote the union of balls of radius $r$ centered at ${\mathcal{D}}$. A function class is simply a set of real-valued functions. Given a set ${\mathcal{D}}$, a function class over ${\mathcal{D}}$ is simply a subset of all real-valued functions on ${\mathcal{D}}$. Given a real number $\mu\geq 0$ and a set ${\mathcal{D}}$, we use ${\mathbf{Q}}_{\mu}$ to denote the class of functions of the form $q_{{\mathbf{x}},{\mathbf{o}}}({\mathbf{y}})=\langle{\mathbf{o}},{\mathbf{y}}-{\mathbf{x}}\rangle+\frac{\mu}{2}\|{\mathbf{y}}-{\mathbf{x}}\|^{2}$, where ${\mathbf{o}}\in{\mathbb{R}}^{d}$ and ${\mathbf{x}}\in{\mathcal{D}}$. A set ${\mathcal{K}}\subseteq{\mathbb{R}}^{d}$ is called a convex set if for all ${\mathbf{x}},{\mathbf{y}}\in{\mathcal{K}}$ and $\alpha\in[0,1]$, we have $\alpha{\mathbf{x}}+(1-\alpha){\mathbf{y}}\in{\mathcal{K}}$. Given a convex set ${\mathcal{K}}$, a function $f:{\mathcal{K}}\to{\mathbb{R}}$ is called convex if, for all ${\mathbf{x}},{\mathbf{y}}\in{\mathcal{K}}$ and $\alpha\in[0,1]$, we have $f(\alpha{\mathbf{x}}+(1-\alpha){\mathbf{y}})\leq\alpha f({\mathbf{x}})+(1-\alpha)f({\mathbf{y}})$. All convex functions are continuous on any point in the relative interior of their domains, but not necessarily at their relative boundary. In this work, we will only focus on continuous functions. If ${\mathbf{x}}\in\operatorname{relint}({\mathcal{K}})$ and $f$ is convex and is differentiable at ${\mathbf{x}}$, then we have $f({\mathbf{y}})-f({\mathbf{x}})\geq\langle\nabla f({\mathbf{x}}),{\mathbf{y}}-{\mathbf{x}}\rangle$, for all ${\mathbf{y}}\in{\mathcal{K}}$. A vector ${\mathbf{o}}\in{\mathbb{R}}^{d}$ is called a subgradient of $f$ at ${\mathbf{x}}$ if $f({\mathbf{y}})-f({\mathbf{x}})\geq\langle{\mathbf{o}},{\mathbf{y}}-{\mathbf{x}}\rangle$, for all ${\mathbf{y}}\in{\mathcal{K}}$. Given a convex set ${\mathcal{K}}$, a function $f:{\mathcal{K}}\to{\mathbb{R}}$ is convex if and only if it has a subgradient at all points ${\mathbf{x}}\in{\mathcal{K}}$. More generally, given $\mu\geq 0$, we say a vector ${\mathbf{o}}\in{\mathbb{R}}^{d}$ is a $\mu$-subgradient of $f$ at ${\mathbf{x}}$ if $f({\mathbf{y}})-f({\mathbf{x}})\geq\langle{\mathbf{o}},{\mathbf{y}}-{\mathbf{x}}\rangle+\frac{\mu}{2}\|{\mathbf{y}}-{\mathbf{x}}\|^{2}$, for all ${\mathbf{y}}\in{\mathcal{K}}$. Given a convex set ${\mathcal{K}}$, a function $f:{\mathcal{K}}\to{\mathbb{R}}$ is $\mu$-strongly convex if and only if it has a $\mu$-subgradient at all points ${\mathbf{x}}\in{\mathcal{K}}$. We use $\nabla^{*}f$ to denote the set of $\mu$-subgradients of $f$.
3 Problem Setup
@@
A query oracle is a function that provides the observation to the agent. Formally, a query oracle for a function $f$ is a map ${\mathcal{Q}}$ defined on ${\mathcal{K}}$ such that for each ${\mathbf{x}}\in{\mathcal{K}}$, the ${\mathcal{Q}}({\mathbf{x}})$ is a random variable taking value in the observation space ${\mathcal{O}}$. The query oracle is called a stochastic value oracle or stochastic zeroth order oracle if ${\mathcal{O}}={\mathbb{R}}$ and $f({\mathbf{x}})={\mathbb{E}}[{\mathcal{Q}}({\mathbf{x}})]$. Similarly, it is called a stochastic (sub)gradient oracle or stochastic first order oracle if ${\mathcal{O}}={\mathbb{R}}^{d}$ and ${\mathbb{E}}[{\mathcal{Q}}({\mathbf{x}})]$ is a (sub)gradient of $f$ at ${\mathbf{x}}$. In all cases, if the random variable takes a single value with probability one, we refer to it as a deterministic oracle. Note that, given a function, there is at most a single deterministic gradient oracle, but there may be many deterministic subgradient oracles. We will use $\nabla$ to denote the deterministic gradient oracle. We say an oracle is bounded by $B$ if its output is always within the Euclidean ball of radius $B$ centered at the origin. We say the agent takes semi-bandit feedback if the oracle is first-order and the agent query function is trivial. Similarly, it takes bandit feedback if the oracle is zeroth-order and the agent query function is trivial [^1] This is a slight generalization of the common use of the term bandit feedback. Usually, bandit feedback refers to the case where the oracle is a deterministic zeroth-order oracle and the agent query function is trivial.. If the agent query function is non-trivial, then we say the agent requires full-information feedback.
@@
An adversary $\operatorname{Adv}$ is a set such that each element ${\mathcal{B}}\in\operatorname{Adv}$, referred to as a realized adversary, is a sequence $({\mathcal{B}}_{1},\cdots,{\mathcal{B}}_{T})$ of functions where each ${\mathcal{B}}_{t}$ maps a tuple $({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{t})\in{\mathcal{K}}^{t}$ to a tuple $(f_{t},{\mathcal{Q}}_{t})$ where $f_{t}\in{\mathbf{F}}$ and ${\mathcal{Q}}_{t}$ is a query oracle for $f_{t}$  [^2] Note that we do not assign a probability to each realized adversary since the notion of regret simply computes the supremum over all realizations.. We say an adversary $\operatorname{Adv}$ is oblivious if for any realization ${\mathcal{B}}=({\mathcal{B}}_{1},\cdots,{\mathcal{B}}_{T})$, all functions ${\mathcal{B}}_{t}$ are constant, i.e., they are independent of $({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{t})$. In this case, a realized adversary may be simply represented by a sequence of functions $(f_{1},\cdots,f_{T})\in{\mathbf{F}}^{T}$ and a sequence of query oracles $({\mathcal{Q}}_{1},\cdots,{\mathcal{Q}}_{T})$ for these functions. We say an adversary is a weakly adaptive adversary if each function ${\mathcal{B}}_{t}$ described above does not depend on ${\mathbf{x}}_{t}$ and therefore may be represented as a map defined on ${\mathcal{K}}^{t-1}$. In this work we also consider adversaries that are fully adaptive, i.e., adversaries with no restriction. Clearly any oblivious adversary is a weakly adaptive adversary and any weakly adaptive adversary is a fully adaptive adversary. Given a function class ${\mathbf{F}}$ and $i\in\{0,1\}$, we use $\operatorname{Adv}^{\text{f}}_{i}({\mathbf{F}})$ to denote the set of all possible realized adversaries with deterministic $i$-th order oracles. If the oracle is instead stochastic and bounded by $B$, we use $\operatorname{Adv}^{\text{f}}_{i}({\mathbf{F}},B)$ to denote such an adversary. Finally, we use $\operatorname{Adv}^{\text{o}}_{i}({\mathbf{F}})$ and $\operatorname{Adv}^{\text{o}}_{i}({\mathbf{F}},B)$ to denote all oblivious realized adversaries with $i$-th order deterministic and stochastic oracles, respectively.

### v4

@@
Abstract.
In this paper, we analyze the problem of online convex optimization in different settings, including different feedback types (full-information/semi-bandit/bandit/etc) in either stochastic or non-stochastic setting and different notions of regret (static adversarial regret/dynamic regret/adaptive regret). This is done through a framework which allows us to systematically propose and analyze meta-algorithms for the various settings described above. We show that any algorithm for online linear optimization with deterministic gradient feedback against fully adaptive adversaries is an algorithm for online convex optimization. We also show that any such algorithm that requires full-information feedback may be transformed to an algorithm with semi-bandit feedback with comparable regret bound. We further show that algorithms that are designed for fully adaptive adversaries using deterministic semi-bandit feedback can obtain similar bounds using only stochastic semi-bandit feedback when facing oblivious adversaries. We use this to describe general meta-algorithms to convert first order algorithms to zeroth order algorithms with comparable regret bounds. Our framework allows us to analyze online optimization in various settings, recovers several results in the literature with a simplified proof technique, and provides new results.
Key words and phrases:
Online Convex Optimization; Meta-algorithms; Regret Minimization; Bandit Feedback; Gradient Estimation
1. Introduction
Online optimization problems represent a class of problems where the decision-maker, referred to as an agent, aims to make sequential decisions in the face of an adversary with incomplete information shalev-shwartz12_onlin_learn_onlin_convex_optim; hazan2016introduction. This setting mirrors a repeated game, where the agent and the adversary engage in a strategic interplay over a finite time horizon, commonly denoted as $T$. The dynamics of this game involve the agent’s decision-making process, the adversary’s strategy, and the information exchange through a query oracle.
@@ 1. Introduction
One of our key contribution lies in establishing that any semi-bandit feedback online linear (or quadratic) optimization algorithm for fully adaptive adversary is also an online convex (or strongly convex) optimization algorithm. While the above result is for semi-bandit feedback, we then show that in online convex optimization for fully adaptive adversary, semi-bandit feedback is generally enough and more information is not needed. Further, we show that algorithms that are designed for fully adaptive adversaries using deterministic semi-bandit feedback can obtain similar bounds using only stochastic semi-bandit feedback when facing oblivious adversaries. Finally, we introduce meta algorithms, based on the variant of classical result of flaxman2005online; shamir17_optim_algor_bandit_zero_order, that convert algorithms for semi-bandit feedback, such as the ones mentioned above, to algorithms for bandit or zeroth-order feedback. In addition to recovering many results in the literature with simplified proofs, we give new static regret guarantees for online strongly convex optimization and adaptive regret guarantees for convex optimization with deterministic zeroth order feedback.
Figure 1. Summary of the main results
@@
(I) We provide a framework that simplifies analysis of meta-algorithms in online learning.
(II) Using this framework, we formulate several meta-algorithms that facilitate the conversion of algorithms across various settings.
(III) In particular, we convert several insights in the literature into precise theorems, for example the idea of using 1-point gradient estimator in online optimization is formalized in Meta-algorithms 2 and 3 and Theorems 10 and 11. Similarly, the insight that “knowing the value of each loss function at two points is almost as useful as knowing the value of each function everywhere” (See agarwal2010optimal) is formalized in Meta-algorithm 4 and Theorem 14.
@@
(i) The first key result in this paper is Theorem 5. This result shows that any algorithm for online linear optimization with deterministic subgradient feedback results in a comparable regret when the function class is replaced by a class of convex functions. This is the key result that allows us to go from online linear optimization to convex optimization in Figure 1 with semi-bandit feedback. We note that this result recovers the result for online convex game in abernethy2008optimal, where the authors analyzed online convex optimization with deterministic gradient feedback and fully adaptive adversary and showed that the min-max regret bounds of online convex optimization is equal to that of online linear optimization (or online quadratic optimization is all convex functions are $\mu$-strongly convex).
@@
(iii) We note that the above results assume that the feedback oracle is deterministic. In the next key result, given in Theorem 9, we relax this assumption. We show that any online optimization algorithm for fully adaptive adversaries using deterministic semi-bandit feedback obtains similar regret bounds when facing an oblivious adversary using only stochastic semi-bandit feedback. Thus, this result allows us to get the results for online convex optimization with stochastic semi-bandit feedback from the results with deterministic semi-bandit feedback, as can be seen in Fig. 1. This result gives Theorem 6.5 in hazan2016introduction as a special case. Further, this allows us to have a result that generalizes both adversarial regret and stochastic regret, as they are commonly defined in the literature.
@@
(iv) The above discussion was mostly focused on full information and semi-bandit feedback settings. However, in many practical setups, we only have access to zeroth order or bandit feedback. In the next key result, given in Theorems 10-11, we use the smoothing trick nemirovsky83_probl_compl_method_effic_optim; flaxman2005online; pedramfar23_unified_approac_maxim_contin_dr_funct, in order to estimate the gradient from the value oracle. Using this estimated gradient, the results for zeroth order and bandit feedback are derived from that for first order and semi-bandit feedback, respectively. These results apply to any optimization problem, not necessarily convex. This approach recovers the results in flaxman2005online with the base algorithm of Online Gradient Descent ($\mathtt{OGD}$). In particular, this approach also matches the SOTA dynamic and adaptive regret bounds for online convex optimization with bandit feedback zhao21_bandit_convex_optim_non_envir; garber22_new_projec_algor_onlin_convex; agarwal2010optimal with a simplified proof.
@@
(v) In the above results, we assumed that we only have access to an unbiased estimate of the zeroth order (or bandit) feedback. In agarwal2010optimal, the setting of multi-point bandit feedback is introduced. Even though their setting and notion of regret is different from our work, one insight is relevant. Their analysis suggests that, in some cases, if we have access to exact values of the functions, knowing the value of each loss function at two points is almost as useful as knowing the value of each function everywhere. We formalize this insight by designing a meta-algorithm based on the two-point gradient estimator agarwal2010optimal; shamir17_optim_algor_bandit_zero_order that allows us to convert algorithms for stochastic first order feedback into algorithms for deterministic zeroth order feedback with the same order of regret bound (Theorem 14). Note that this result holds for convex and non-convex optimization as it only relies on the function class being Lipschitz. This result is shown to recover the result for dynamic regret for Improved Ader zhao21_bandit_convex_optim_non_envir, while achieving two new results for adaptive regret for convex function of $O(\sqrt{T})$ and static regret for strongly convex function of $O(\log T)$.
@@
Our work sheds light on the relation between different problems, algorithms and techniques used in online optimization and provides general meta-algorithms that allow conversion between different problem settings. For instance, we can use the results for a deterministic algorithm for online linear optimization with oblivious adversaries and deterministic semi-bandit feedback to that for a stochastic algorithm for online convex optimization with oblivious adversaries and stochastic bandit feedback using the different arrows as in Fig. 1. The key technical novelty is the recognition that many results, previously understood only in very limited contexts, actually apply more broadly. This broader applicability not only simplifies the analysis of numerous existing results within the field but also enables the derivation of multiple new results. For instance, results in abernethy2008optimal; hazan2016introduction; flaxman2005online; zhao21_bandit_convex_optim_non_envir; agarwal2010optimal; shamir17_optim_algor_bandit_zero_order as mentioned above can be demonstrated as corollaries of these more generalized results. Further, we provide new results for deterministic zeroth-order feedback case, showing that adaptive regret for convex function is $O(\sqrt{T})$ and static regret for strongly convex function is $O(\log T)$.
@@
Related works: The term “online convex optimization” was first defined in zinkevich03_onlin, where Online Gradient Descent ($\mathtt{OGD}$) was proposed, while this setting was first introduced in gordon99_regret. The Follow-The-Perturbed-Leader ($\mathtt{FTPL}$) for online linear optimization was introduced by kalai05_effic, inspired by Follow-The-Leader algorithm originally suggested by hannan57. Follow-The-Regularized-Leader ($\mathtt{FTRL}$) in the context of online convex optimization was coined in shalev-shwartz07; shalev-shwartz07_onlin. The equivalence of $\mathtt{FTRL}$ and Online Mirror Descent ($\mathtt{OMD}$) was observed by hazan10_extrac. Non-stationary measures of regret have been studied in zinkevich03_onlin; hazan09_effic; besbes15_non_station_stoch_optim; daniely15_stron_adapt_onlin_learn; zhang18_dynam_regret_stron_adapt_method; zhang18_adapt_onlin_learn_dynam_envir; zhao20_dynam_regret_convex_smoot_funct; zhao21_bandit_convex_optim_non_envir; lu23_projec_adapt_regret_member_oracl; wang24_non_projec_free_onlin_learn; garber22_new_projec_algor_onlin_convex and many other papers. We refer the reader to surveys shalev-shwartz12_onlin_learn_onlin_convex_optim and hazan2016introduction for more details.
2. Background and Notations
@@
For a set ${\mathcal{S}}\subseteq{\mathbb{R}}^{d}$, we define its affine hull $\operatorname{aff}({\mathcal{S}})$ to be the set of all points of the form $\sum_{i=1}^{m}\alpha_{i}{\mathbf{x}}_{i}$ for integer $m\geq 1$, points $({\mathbf{x}}_{i})_{i=1}^{m}$ in $\mathcal{S}$}, and real numbers $(\alpha_{i})_{i=1}^{m}$. The relative interior of $\mathcal{S}$} is defined as
$\operatorname{relint}({\mathcal{S}}):=\{{\mathbf{x}}\in{\mathcal{S}}\mid\exists r>0,{\mathbb{B}}_{r}({\mathbf{x}})\cap\operatorname{aff}({\mathcal{S}})\subseteq{\mathcal{S}}\}.$
@@
We use ${\mathbb{B}}_{r}({\mathcal{S}})$ to denote the union of balls of radius $r$ centered at $\mathcal{S}$} and use $\operatorname{diam}({\mathcal{S}}):=\sup_{{\mathbf{x}},{\mathbf{y}}\in{\mathcal{S}}}\|{\mathbf{x}}-{\mathbf{y}}\|$ to denote the diameter of $\mathcal{S}$}. A function class is simply a set of real-valued functions. Given a set $\mathcal{S}$}, a function class over $\mathcal{S}$} is simply a subset of all real-valued functions on $\mathcal{S}$}. Given a real number $\mu\geq 0$ and a set $\mathcal{S}$}, we use ${\mathbf{Q}}_{\mu}$ to denote the class of functions of the form $q_{{\mathbf{x}},{\mathbf{o}}}({\mathbf{y}})=\langle{\mathbf{o}},{\mathbf{y}}-{\mathbf{x}}\rangle+\frac{\mu}{2}\|{\mathbf{y}}-{\mathbf{x}}\|^{2}$, where ${\mathbf{o}}\in{\mathbb{R}}^{d}$ and ${\mathbf{x}}\in{\mathcal{S}}$. A set ${\mathcal{K}}\subseteq{\mathbb{R}}^{d}$ is called a convex set if for all ${\mathbf{x}},{\mathbf{y}}\in{\mathcal{K}}$ and $\alpha\in[0,1]$, we have $\alpha{\mathbf{x}}+(1-\alpha){\mathbf{y}}\in{\mathcal{K}}$. Given a convex set $\mathcal{K}$}, a function $f:{\mathcal{K}}\to{\mathbb{R}}$ is called convex if, for all ${\mathbf{x}},{\mathbf{y}}\in{\mathcal{K}}$ and $\alpha\in[0,1]$, we have $f(\alpha{\mathbf{x}}+(1-\alpha){\mathbf{y}})\leq\alpha f({\mathbf{x}})+(1-\alpha)f({\mathbf{y}})$. All convex functions are continuous on any point in the relative interior of their domains, but not necessarily at their relative boundary. In this work, we will only focus on continuous functions. If ${\mathbf{x}}\in\operatorname{relint}({\mathcal{K}})$ and $f$ is convex and is differentiable at $\mathbf{x}$}, then we have $f({\mathbf{y}})-f({\mathbf{x}})\geq\langle\nabla f({\mathbf{x}}),{\mathbf{y}}-{\mathbf{x}}\rangle$, for all ${\mathbf{y}}\in{\mathcal{K}}$. A vector ${\mathbf{o}}\in{\mathbb{R}}^{d}$ is called a subgradient of $f$ at $\mathbf{x}$} if $f({\mathbf{y}})-f({\mathbf{x}})\geq\langle{\mathbf{o}},{\mathbf{y}}-{\mathbf{x}}\rangle$, for all ${\mathbf{y}}\in{\mathcal{K}}$. More generally, given $\mu\geq 0$, we say a vector ${\mathbf{o}}\in{\mathbb{R}}^{d}$ is a $\mu$-subgradient of $f$ at $\mathbf{x}$} if $f({\mathbf{y}})-f({\mathbf{x}})\geq\langle{\mathbf{o}},{\mathbf{y}}-{\mathbf{x}}\rangle+\frac{\mu}{2}\|{\mathbf{y}}-{\mathbf{x}}\|^{2}$, for all ${\mathbf{y}}\in{\mathcal{K}}$. Given a convex set $\mathcal{K}$}, a Lipschitz continuous function $f:{\mathcal{K}}\to{\mathbb{R}}$ is convex (resp. $\mu$-strongly convex) if and only if it has a ($\mu$-)subgradient at all points ${\mathbf{x}}\in{\mathcal{K}}$. We use $\nabla^{*}f$ to denote the set of $\mu$-subgradients of $f$.
3. Problem Setup
@@
A query oracle is a function that provides the observation to the agent. Formally, a query oracle for a function $f$ is a map $\mathcal{Q}$} defined on $\mathcal{K}$} such that for each ${\mathbf{x}}\in{\mathcal{K}}$, the ${\mathcal{Q}}({\mathbf{x}})$ is a random variable taking value in the observation space $\mathcal{O}$}. The query oracle is called a stochastic value oracle or stochastic zeroth order oracle if $\mathcal{O}$$\mathbb{R}$} and $f({\mathbf{x}})={\mathbb{E}}[{\mathcal{Q}}({\mathbf{x}})]$. Similarly, it is called a stochastic (sub)gradient oracle or stochastic first order oracle if ${\mathcal{O}}={\mathbb{R}}^{d}$ and ${\mathbb{E}}[{\mathcal{Q}}({\mathbf{x}})]$ is a (sub)gradient of $f$ at $\mathbf{x}$}. In all cases, if the random variable takes a single value with probability one, we refer to it as a deterministic oracle. Note that, given a function, there is at most a single deterministic gradient oracle, but there may be many deterministic subgradient oracles. We will use $\nabla$ to denote the deterministic gradient oracle. We say an oracle is bounded by $B$ if its output is always within the Euclidean ball of radius $B$ centered at the origin. We say the agent takes semi-bandit feedback if the oracle is first-order and the agent query function is trivial. Similarly, it takes bandit feedback if the oracle is zeroth-order and the agent query function is trivial  [^1] This is a slight generalization of the common use of the term bandit feedback. Usually, bandit feedback refers to the case where the oracle is a deterministic zeroth-order oracle and the agent query function is trivial, while our definition allows for stochastic oracles.. If the agent query function is non-trivial, then we say the agent requires full-information feedback.
@@
An adversary $\operatorname{Adv}$ is a set such that each element ${\mathcal{B}}\in\operatorname{Adv}$, referred to as a realized adversary, is a sequence $({\mathcal{B}}_{1},\cdots,{\mathcal{B}}_{T})$ of functions where each ${\mathcal{B}}_{t}$ maps a tuple $({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{t})\in{\mathcal{K}}^{t}$ to a tuple $(f_{t},{\mathcal{Q}}_{t})$ where $f_{t}\in{\mathbf{F}}$ and ${\mathcal{Q}}_{t}$ is a query oracle for $f_{t}$  [^2] Note that we do not assign a probability to each realized adversary since the notion of regret simply computes the supremum over all realizations.. We say an adversary $\operatorname{Adv}$ is oblivious if for any realization ${\mathcal{B}}=({\mathcal{B}}_{1},\cdots,{\mathcal{B}}_{T})$, all functions ${\mathcal{B}}_{t}$ are constant, i.e., they are independent of $({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{t})$. In this case, a realized adversary may be simply represented by a sequence of functions $(f_{1},\cdots,f_{T})\in{\mathbf{F}}^{T}$ and a sequence of query oracles $({\mathcal{Q}}_{1},\cdots,{\mathcal{Q}}_{T})$ for these functions. In this work we also consider adversaries that are fully adaptive, i.e., adversaries with no restriction.  [^3] Another form of adversary considered in literature is a weakly adaptive adversary where each function ${\mathcal{B}}_{t}$ described above does not depend on ${\mathbf{x}}_{t}$ and therefore may be represented as a map defined on ${\mathcal{K}}^{t-1}$. Clearly any oblivious adversary is a weakly adaptive adversary and any weakly adaptive adversary is a fully adaptive adversary. Given a function class $\mathbf{F}$} and $i\in\{0,1\}$, we use $\operatorname{Adv}^{\text{f}}_{i}({\mathbf{F}})$ to denote the set of all possible realized adversaries with deterministic $i$-th order oracles. If the oracle is instead stochastic and bounded by $B$, we use $\operatorname{Adv}^{\text{f}}_{i}({\mathbf{F}},B)$ to denote such an adversary. Finally, we use $\operatorname{Adv}^{\text{o}}_{i}({\mathbf{F}})$ and $\operatorname{Adv}^{\text{o}}_{i}({\mathbf{F}},B)$ to denote all oblivious realized adversaries with $i$-th order deterministic and stochastic oracles, respectively.